• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name vasulk

    Log entry time 18:04:00 on February14,2003

    Entry number 96484

    keyword=Target Status

    Today the target and field survey were completed at 15:00. After the target survey, Exodus was removed fully intact from the scattering chamber. Preliminary results from the target survey looked good, however there was still about a 2 mm offset in the vertical direction due to expansion from the oven. Repeatability tests looked very good.

    The field survey for 0 and 180 degrees were withing 0.5 degrees, with 0 degrees almost being perfect. However 270 degrees showed ~ 2 degree offset. Unfortunately we did not have the adequate amount of time to let the large coils warm up due to time constraints. It is possible once the coils warmed up, the offset would have been smaller. The initial survey, in December, had 270 degrees within 0.5 degrees.

    After the survey, JP and I removed most of the scattering chamber covers in preparation for water calibrations this weekend.

    17:15 Holding field initialized for water calibrations